# System-2 Reasoning via Generality and Adaptation

### v1

@@ 4.4 Reinforcement Learning Strengthens Adaptation and Multi-Step Logical Reasoning
System-2 reasoning requires not only logical deduction but also adaptability in dynamic environments and the ability to execute multi-step reasoning. Reinforcement Learning (RL) is particularly suited for this, as it enables AI systems to learn through trial and error by optimizing actions based on rewards received from the environment. By continuously adjusting actions to maximize long-term rewards, RL enhances AI’s ability to adapt to unfamiliar tasks and perform multi-step reasoning, which are both critical for System-2 reasoning.
@@
In ARC tasks, RL agents can be trained on ARCLE [34] and leveraged to learn optimal sequences of transformations. For instance, HRL can decompose a complex ARC task into sub-tasks, where each layer of the hierarchy handles a different level of abstraction, simplifying the reasoning process. MBRL can simulate the outcomes of different transformations, enabling AI to plan and choose the most effective sequence of actions [35]. RRL, by focusing on object relationships, allows AI to reason about how different transformations affect spatial arrangements, adapting to new tasks that involve object interactions. Through these methods, RL strengthens AI’s ability to perform multi-step logical reasoning and adapt to novel environments, bringing it closer to System-2 reasoning.
@@ 5 Conclusion
This paper explored the limitations of existing AI approaches, such as program synthesis, transformers, and large language models (LLMs), in solving ARC tasks, particularly in the context of generality and adaptation. While these methods have demonstrated notable improvements in generality, they consistently struggle with adaptation to novel and unfamiliar environments. Their reliance on predefined structures, patterns, or vast pre-training makes it challenging for current AI systems to achieve the deeper, more flexible reasoning associated with System-2 capabilities.
To address these challenges, we proposed four key research directions aimed at enhancing both generality and adaptation for System-2 reasoning: (1) learning human intentions from action sequences, (2) integrating symbolic and neural hybrid models, (3) employing meta-learning for adaptation, and (4) using reinforcement learning for multi-step logical reasoning. Each of these approaches targets specific limitations of current methods, helping models better handle abstract reasoning, generalize across various domains, and dynamically adapt to novel tasks.
@@
Advancing towards System-2 reasoning will require these approaches to work in synergy, allowing models to generalize across diverse tasks and adapt seamlessly to unforeseen scenarios. By pursuing these research directions, AI systems can evolve closer to AGI-level reasoning, demonstrating not only pattern recognition but true human-level reasoning, logical deduction, and abstract thinking.

### v3

@@ 4.4 Reinforcement Learning Strengthens Adaptation and Multi-Step Logical Reasoning
System 2 reasoning requires not only logical deduction but also adaptability in dynamic environments and the ability to execute multi-step reasoning. Reinforcement Learning (RL) is particularly suited for this, as it enables AI systems to learn through trial and error by optimizing actions based on rewards received from the environment. By continuously adjusting actions to maximize long-term rewards, RL enhances AI’s ability to adapt to unfamiliar tasks and perform multi-step reasoning, which are both critical for System 2 reasoning.
@@
In ARC tasks, RL agents can be trained on ARCLE [34] and leveraged to learn optimal sequences of transformations. For instance, HRL can decompose a complex ARC task into sub-tasks, where each layer of the hierarchy handles a different level of abstraction, simplifying the reasoning process. MBRL can simulate the outcomes of different transformations, enabling AI to plan and choose the most effective sequence of actions [35]. RRL, by focusing on object relationships, allows AI to reason about how different transformations affect spatial arrangements, adapting to new tasks that involve object interactions. Through these methods, RL strengthens AI’s ability to perform multi-step logical reasoning and adapt to novel environments, bringing it closer to System 2 reasoning.
@@ 5 Conclusion
This paper explored the limitations of existing AI approaches, such as program synthesis, transformers, and large language models (LLMs), in solving ARC tasks, particularly in the context of generality and adaptation. While these methods have demonstrated notable improvements in generality, they consistently struggle with adaptation to novel and unfamiliar environments. Their reliance on predefined structures, patterns, or vast pre-training makes it challenging for current AI systems to achieve the deeper, more flexible reasoning associated with System 2 capabilities.
To address these challenges, we proposed four key research directions aimed at enhancing both generality and adaptation for System 2 reasoning: (1) learning human intentions from action sequences, (2) integrating symbolic and neural hybrid models, (3) employing meta-learning for adaptation, and (4) using reinforcement learning for multi-step logical reasoning. Each of these approaches targets specific limitations of current methods, helping models better handle abstract reasoning, generalize across various domains, and dynamically adapt to novel tasks.
@@
Advancing towards System 2 reasoning will require these approaches to work in synergy, allowing models to generalize across diverse tasks and adapt seamlessly to unforeseen scenarios. By pursuing these research directions, AI systems can evolve closer to AGI-level reasoning, demonstrating not only pattern recognition but true human-level reasoning, logical deduction, and abstract thinking.